OpenStudy (anonymous):

Find the point of intersection, X, of the perpendicular bisectors, y=-x+4 and y=x-1

OpenStudy (anonymous):

-x+4=x-1 2 x =5 x=5/2

OpenStudy (anonymous):

y=5/2 -1=5/2 -2/2=3/2

OpenStudy (anonymous):

(5/2,3/2) is the answer

OpenStudy (anonymous):

thank you! but how do you work it out? i have an assignment due tomorrow and im not sure how to work out the equation